# ReCalKV: Low-Rank KV Cache Compression via Head Reordering and Offline Calibration

### v2

@@ 6 Conclusion
In this work, we prpose ReCalKV, a post-training KV cache compression framework tailored for efficient long-context reasoning in LLMs. By exploiting the distinct characteristics of Keys and Values in the attention mechanism, ReCalKV applies Head-wise Similarity-aware Reordering (HSR) and grouped SVD to compress Keys, while employing Offline Calibration and Matrix Fusion (OCMF) to compress Values. This design reduces hidden dimensions with minimal additional computation and preserves model performance under high compression ratios. Experimental results demonstrate that ReCalKV consistently outperforms existing low-rank compression methods, offering a practical and effective solution for memory-efficient LLM serving. Moreover, it can be combined with quantization to achieve higher compression with minimal performance loss. We believe this work provides a promising direction for scalable, long-context LLM deployment.

### v3

@@ 5 Conclusion
In this work, we propose ReCalKV, a post-training KV cache compression framework tailored for efficient long-context reasoning in LLMs. By exploiting the distinct characteristics of Keys and Values in the attention mechanism, ReCalKV applies Head-wise Similarity-aware Reordering (HSR) and grouped SVD to compress Keys, while employing Offline Value Calibration (OVC) to compress Values. This design reduces hidden dimensions with minimal additional computation and preserves model performance under high compression ratios. Experimental results demonstrate that ReCalKV consistently outperforms existing low-rank compression methods, offering a practical and effective solution for memory-efficient LLM serving. Moreover, it can be combined with quantization to achieve higher compression with minimal performance loss. This work offers a promising direction for scalable and efficient deployment of long-context LLMs.